OpenStudy (vortish):

solve the equation (x-1)(2x+9)= 0

OpenStudy (anonymous):

Set each of the factors equal to zero: x-1=0 2x+9=0

OpenStudy (vortish):

x=1 x=-3

OpenStudy (anonymous):

Check the second one 2(-3)+9 = 3 =/= 0. It's not correct

OpenStudy (vortish):

x=-1 2x=-9

OpenStudy (vortish):

have to find the value for x for the seocnd one

OpenStudy (vortish):

x={4.5,1}